OpenStudy (abbycross167):

Can someone please help me answer this math question? Please? I'll fan and give a medal!! Refer to Example 21 Question - When a diver jumps from a ledge, the vertical component of his motion can e modeled by the vertical motion model: h = -16t^2 + 8t + 48. Use factoring to determine how many seconds it takes for the diver to enter the water. Show the steps as outlined in the lesson content.

OpenStudy (abbycross167):

Example 21 - Sam dropped a ball from the top of a building 100 feet high. The height of the ball is represented by the equation h = -t2 + 100 where t represents the time in seconds. Use factoring to determine how many seconds it takes for the ball to reach the ground. Solution: Since the height of the ground is 0, let h = 0. The equation becomes 0 = -t^2 + 100 or -t^2 + 100 = 0. Factoring: -1(t^2 – 100) = 0 Factoring: -1 (t+10)(t-10) = 0 Using the Zero Product Property either factor has to equal 0: t+ 10 = 0 or t-10 = 0 t = -10 or t = 10 t = -10 has no meaning in this problem, but t = 10 does. So the ball will hit the ground in 10 seconds.

OpenStudy (greatlife44):

the only forces acting on this person are the force of gravity which pulls him downward. we know that once he hits the water h =0 the vertical height that is.  '